# Accurate Network Traffic Matrix Prediction via LEAD: an LLM-Enhanced Adapter-Based Conditional Diffusion Model

### v1

@@ Abstract
Driven by the evolution toward 6G and AI-native edge intelligence, network operations increasingly require predictive and risk-aware adaptation under stringent computation and latency constraints. Network Traffic Matrix (TM), which characterizes flow volumes between nodes, is a fundamental signal for proactive traffic engineering. However, accurate TM forecasting remains challenging due to the stochastic, non-linear, and bursty nature of network dynamics. Existing discriminative models often suffer from over-smoothing and provide limited uncertainty awareness, leading to poor fidelity under extreme bursts. To address these limitations, we propose LEAD, a Large Language Model (LLM)-Enhanced Adapter-based conditional Diffusion model. First, LEAD adopts a “Traffic-to-Image” paradigm to transform traffic matrices into RGB images, enabling global dependency modeling via vision backbones. Then, we design a “Frozen LLM with Trainable Adapter” model, which efficiently captures temporal semantics with limited computational cost. Moreover, we propose a Dual-Conditioning Strategy to precisely guide a diffusion model to generate complex, dynamic network traffic matrices. Experiments on the Abilene and GÉANT datasets demonstrate that LEAD outperforms all baselines. On the Abilene dataset, LEAD attains a remarkable 45.2% reduction in RMSE against the best baseline, with the error margin rising only marginally from 0.1098 at one-step to 0.1134 at 20-step predictions. Meanwhile, on the GÉANT dataset, LEAD achieves a 0.0258 RMSE at 20-step prediction horizon which is 27.3% lower than the best baseline.
Index Terms:
Network Traffic Matrix Prediction, Foundation Models, Large Language Models, Diffusion Models, AI-Native Edge Networks.
1 Introduction
@@
As wireless systems evolve toward 6G and AI-native edge intelligence, network operations increasingly require predictive adaptation under stringent latency and resource constraints across distributed edge nodes. Meanwhile, the explosive growth of Internet traffic, driven by emerging applications like ultra-high-definition video streaming, cloud computing, and the Internet of Things (IoT), continues to place unprecedented demands on modern communication networks [1], [2]. Central to effective network management is the network Traffic Matrix (TM), which captures the time-varying aggregate flow volume exchanged between network nodes over a given time interval. By offering a global view of network demand, TMs serve as the cornerstone for critical tasks such as Traffic Engineering (TE), Capacity Planning, and Service Level Agreement (SLA) assurance [3], [4], [5]. However, obtaining real-time TMs is often plagued by high measurement overhead and significant transmission delays, making reactive adjustments insufficient for guaranteeing Quality of Service (QoS) in modern high-speed networks. Consequently, TM prediction, defined as forecasting network traffic loads based on historical measurements, has emerged as an indispensable capability because it empowers network operators to proactively optimize routing protocols and mitigate congestion before it physically occurs, shifting network management from a reactive to a proactive paradigm.
@@
Traditional TM prediction often uses statistical time-series and regression models such as ARIMA and Kalman filtering [6]. While efficient, they typically rely on simplified assumptions and weakly capture network-wide spatial coupling and multi-timescale temporal dynamics, which can reduce robustness under bursty traffic in large backbone networks.
@@
Deep learning methods provide a natural remedy by learning richer temporal dependencies from historical measurements. Recurrent Neural Networks (RNNs) improve sequential modeling but may still miss spatial interactions and long-range effects. Methods based on Graph Neural Network (GNN) address this by encoding adjacency structure, and recent work further explores spatio-temporal attention for TM modeling [7]. However, they can be sensitive to graph specification and over-smoothing, weakening responsiveness to peaks and anomalies. Moreover, most existing predictors are deterministic and provide only point forecasts, thus failing to quantify uncertainty and limiting their utility for risk-aware operation under strict SLAs.
@@
We therefore hypothesize that Large Language Models (LLMs) can provide semantic abstraction and reasoning to strengthen conditioning for generative TM forecasting. However, directly fine-tuning massive LLMs for a niche network-traffic task is often computationally prohibitive, so we adopt parameter-efficient adaptation inspired by the LLaMA-Adapter [12], which enable the efficient alignment of frozen LLMs with the network traffic domain via lightweight adaptation.
@@
  • We propose LEAD, a novel model that orchestrates a deep synergy between LLMs and conditional diffusion models for network traffic prediction. Beyond simple integration, we establish a dependency where the LLM acts as a semantic reasoner to extract high-level temporal dynamics. This logic serves as explicit guidance for the diffusion model, ensuring the generated traffic matrices are not only numerically precise but also consistent with realistic network behaviors and topological constraints, while keeping the LLM deployable via parameter-efficient adaptation.
@@
The remainder of this paper is organized as follows: Section 2 reviews related works. Section 3 details the methodology of the proposed LEAD model. Section 4 reports our experimental results, including ablation studies. Section 5 concludes the paper and proposes future research directions.
2 Related Works
2.1 Network Traffic Prediction Methods
@@
Early methodologies rely heavily on statistical techniques. Historical Average (HA), ARIMA [13], and Kalman filtering [6] are classic examples that model linear dependencies, while Vector Auto-Regression (VAR) [14] extends these to multivariate settings. Despite their efficiency, these models are limited by stationarity assumptions and often struggle to capture the non-linear dynamics of complex networks.
To model non-linear temporal correlations, RNNs, such as LSTM and GRU [15], are introduced and become foundational baselines. However, these methods treat network traffic streams as isolated sequences, largely neglecting the underlying topological structure and spatial interactions between nodes.
Consequently, Spatio-Temporal Graph Neural Networks (STGNNs) have been widely adopted to jointly model spatial and temporal dependencies. Representative works like DCRNN [16] and STGCN [17] combine graph convolutions with recurrent or convolutional units to capture directed spatial flows. To mitigate the reliance on predefined graph structures, methods such as Graph WaveNet [18] and AGCRN [19] propose adaptive graph learning strategies to infer latent correlations automatically. Furthermore, recent studies in the database community have actively explored dynamic and multiscale structure learning to capture evolving dependencies. For instance, METRO [20] introduces a generic framework leveraging multiscale temporal graphs to model variable correlations, while Zhao et al. [21] propose explicitly generating discrete graph structures to adapt to changing time-series relationships. In parallel, researchers have explored advanced representation learning paradigms. For instance, ST-SSL [22] leverages self-supervised learning schemes to enhance robustness against data heterogeneity, while MC-STL [23] employs multi-channel architectures to disentangle complex trend and seasonality components. More recently, researchers have explored incorporating domain knowledge into STGNNs through attention mechanisms [7].
@@ 2.3 LLMs
In time-series analysis, for instance, researchers have creatively adapted LLMs by tokenizing numerical sequences. Zhou et al. [30] treats time-series as pseudo-sentences for zero-shot forecasting, while frameworks like Time-LLM [31] and UniTime [32] have demonstrated that a frozen LLM’s internal reasoning capacity can be harnessed for numerical prediction through clever input reprogramming and prompting, effectively bypassing the need for full fine-tuning.
@@ 3 Methodology
    Encoding: Historical traffic matrices $\textbf{X}_{1:T_{in}}$ are encoded into spatio-temporal visual tokens, preserving high-frequency details.
@@
    Generation: These conditions guide a U-Net based Diffusion Model to iteratively denoise Gaussian noise, reconstructing the future TMs $\hat{\textbf{X}}_{{T_{in}+1}:{T_{in}+T_{out}}}$.
3.1 Data Representation
@@
The transformation’s principal advantage lies in its creation of a representational bridge to CNNs. By transposing network data into the visual domain, we leverage the intrinsic capabilities of vision models to identify structurally analogous patterns: intensity gradients become indicators of high-traffic links, while contiguous dark regions reveal network congestion clusters. This approach not only maintains the topological fidelity of the original data but also unlocks sophisticated feature extraction methodologies traditionally applied to visual analysis, thereby facilitating the discovery of spatially distributed patterns within network traffic flows.
3.2 Spatio-Temporal Perception via Detail-Preserving Encoding
@@
To aggregate the sequence of visual features $\textbf{H}_{v}^{1:T}$ into a coherent spatio-temporal representation, we introduce an enhanced temporal aggregator. Unlike bidirectional methods including Bi-LSTM, traffic prediction must strictly adhere to causality. We employ a Transformer-based aggregator with a causal mask, ensuring that the representation at step $t$ is based solely on the previous steps.
@@
3.3 LLM-Based Reasoning with Adapter
@@
3.4 Dual-Conditioning Strategy
@@
3.5 Conditional Diffusion Generation
@@
As depicted in Fig. 4, we introduce an Adaptive Group Normalization (AdaGN) mechanism to inject the global context $c_{global}$ into the U-Net backbone. AdaGN dynamically modulates the U-Net’s intermediate feature maps by learning scale and shift parameters derived directly from $c_{global}$, thereby guiding the denoising trajectory with global semantic constraints: $\mathrm{AdaGN}(\mathbf{f},c_{global})=\gamma_{s}(c_{global})\cdot\mathrm{GroupNorm}(\mathbf{f})+\beta_{s}(c_{global})$. Here, $\gamma_{s}$ and $\beta_{s}$ are scale and shift parameters predicted from $c_{global}$ via a shallow MLP. This mechanism effectively shifts the distribution of the generated traffic to match the predicted global congestion level. The sequential condition $c_{seq}$ acts as the Key and Value in the cross-attention layers of the U-Net, while the intermediate feature maps of the U-Net serve as Queries.
@@
During inference, we employ the Denoising Diffusion Implicit Models (DDIM) sampler to accelerate the generation process, allowing us to produce the prediction $\hat{\textbf{X}}_{T+1}$ in fewer steps than the 1000-step denoising process.
@@ 4.2 Baselines and Metrics
  • DCRNN & STGCN: DCRNN (Diffusion Convolutional Recurrent Neural Network) integrates diffusion convolution with GRU, while STGCN (Spatial-Temporal Graph Convolutional Network) employs graph convolutions and 1D temporal convolutions for faster inference. They both rely on predefined adjacency matrices to capture spatial dependencies [16], [17].
@@
  • ST-SSL & MC-STL: ST-SSL (Spatio-Temporal Self-Supervised Learning) employs self-supervised learning with pretext tasks to learn robust patterns resilient to noise. MC-STL (Mask- and Contrast-enhanced Spatio-Temporal Learning) decomposes traffic data into multiple channels to model distinct temporal dynamics separately [22], [23].
@@
  • ViT-LSTM: This method is particularly relevant as it shares the “Traffic-to-Image” paradigm with our approach. It converts traffic matrices into image sequences and utilizes ViT to extract spatial features, followed by LSTM for temporal modeling. By comparing with ViT-LSTM, we can directly validate the superiority of our generative diffusion backbone over traditional discriminative visual models [8].
The primary metrics are pixel-wise root mean squared error (RMSE) and mean absolute error (MAE), computed on normalized color value of the first channel of RGB images. Given ground truth $X\in\mathbb{R}^{N\times N}$ and prediction $\hat{X}\in\mathbb{R}^{N\times N}$, we have:
@@
4.3 Setup and Implementation Details
@@
To ensure a rigorous evaluation without data leakage, we adopt a strict chronological splitting strategy for both Abilene and GÉANT datasets. The data is divided into training, validation, and testing sets with a ratio of 7:1.5:1.5. Moreover, we employ a sliding window mechanism to generate samples. The model utilizes the historical traffic sequence of length $\bm{T_{in}=24}$ to predict the TMs at the next $\bm{T_{out}}$ time steps.
@@
  • Conditioning: The model integrates global context via FiLM and local details via cross-attention, with a diffusion noise schedule of 1000 steps.
@@
TABLE I: Performance comparison of different models on Abilene dataset.
@@
TABLE II: Performance comparison of different models on GÉANT dataset.
@@
4.4 Main Results and Analysis
The quantitative comparison between our proposed LEAD and nine baselines on the Abilene and GÉANT datasets is presented in Table  II and Table  II respectively, covering prediction horizon from short-term ($T_{out}=1$) to long-term ($T_{out}=20$). Overall, LEAD achieves outstanding performance across both datasets. To provide a more intuitive understanding of the performance divergence across different paradigms, we contrast our model with STGNNs in Fig. 5a-Fig. 5b and with multi-scale methods in Fig. 5c-Fig. LABEL:fig:.

### v2

# Accurate Network Traffic Matrix Prediction via LEAD: a Large Language Model-Enhanced Adapter-Based Conditional Diffusion Model

@@ Abstract
In the era of AI-native 6G, network operations are evolving from reactive, rule-based paradigms to proactive, cognitive frameworks. To realize this vision, accurate Network Traffic Matrix (TM) prediction serves as a critical sensing mechanism, providing the situational awareness necessary for intelligent resource orchestration. However, accurate TM forecasting remains challenging due to the stochastic, non-linear, and bursty nature of network dynamics. Existing discriminative models often suffer from over-smoothing and provide limited uncertainty awareness, leading to poor fidelity under extreme bursts. To address these limitations, we propose LEAD, a Large Language Model-Enhanced Adapter-based conditional Diffusion model. First, LEAD adopts a “Traffic-to-Image” paradigm to transform traffic matrices into RGB images, enabling global dependency modeling via vision backbones. Then, we design a “Frozen LLM with Trainable Adapter” model, which efficiently captures temporal semantics with limited computational cost. Moreover, we propose a Dual-Conditioning Strategy to precisely guide a diffusion model to generate complex, dynamic network traffic matrices. Experiments on the Abilene and GÉANT datasets demonstrate that LEAD outperforms all baselines. On the Abilene dataset, LEAD attains a remarkable 45.2% reduction in RMSE against the best baseline, with the error margin rising only marginally from 0.1098 at one-step to 0.1134 at 20-step predictions. Meanwhile, on the GÉANT dataset, LEAD achieves a 0.0258 RMSE at 20-step prediction horizon which is 27.3% lower than the best baseline.
I Introduction
@@
As wireless systems evolve toward 6G communication, network operations increasingly require cognitive, closed-loop predictive adaptation under stringent latency and resource constraints across distributed edge nodes. Meanwhile, the explosive growth of Internet traffic, driven by emerging applications like ultra-high-definition video streaming, cloud computing, and the Internet of Things (IoT), continues to place unprecedented demands on modern communication networks [1], [2], [3]. Central to effective network management is the network Traffic Matrix (TM), which captures the time-varying aggregate flow volume exchanged between network nodes over a given time interval. By offering a global view of network demand, TMs serve as the cornerstone for critical tasks such as Traffic Engineering (TE), Capacity Planning, and Service Level Agreement (SLA) assurance [4], [5], [6]. In these operational loops, accurately anticipating bursts and traffic volatility is critical for risk-aware actions rather than purely reactive responses. However, obtaining real-time TMs is often plagued by high measurement overhead and significant transmission delays, making reactive adjustments insufficient for guaranteeing Quality of Service (QoS) in modern high-speed networks. Consequently, TM prediction, defined as forecasting network traffic loads based on historical measurements, has emerged as an indispensable capability because it empowers network operators to proactively optimize routing protocols and mitigate congestion before it physically occurs, shifting network management from a reactive to a proactive paradigm.
@@
Traditional TM prediction often uses statistical time-series and regression models such as ARIMA (Autoregressive Integrated Moving Average) and Kalman filtering [7]. While efficient, they typically rely on simplified assumptions and weakly capture network-wide spatial coupling and multi-timescale temporal dynamics, which can reduce robustness under bursty traffic in large backbone networks.
@@
Deep learning methods provide a natural remedy by learning richer temporal dependencies from historical measurements. RNNs (Recurrent Neural Networks) improve sequential modeling but may still miss spatial interactions and long-range effects. Methods based on GNN (Graph Neural Network) address this by encoding adjacency structure, and recent work further explores spatio-temporal attention for TM modeling [8]. However, they can be sensitive to graph specification and over-smoothing, weakening responsiveness to peaks and anomalies. Moreover, most existing predictors are deterministic and provide only point forecasts, thus failing to quantify uncertainty and limiting their utility for risk-aware operation under strict SLAs.
@@
We therefore hypothesize that Large Language Models (LLMs) can provide semantic abstraction and reasoning to strengthen conditioning for generative TM forecasting. However, directly fine-tuning massive LLMs for a niche network-traffic task is often computationally prohibitive, so we adopt parameter-efficient adaptation inspired by the LLaMA-Adapter [13], which enables the efficient alignment of frozen LLMs with the network traffic domain via lightweight adaptation.
@@
  • We propose LEAD, a novel model that orchestrates a deep synergy between LLMs and conditional diffusion models for network traffic prediction. Beyond simple integration, we establish a dependency where the LLM acts as a semantic reasoner to extract high-level temporal dynamics. This logic serves as explicit guidance for the diffusion model, ensuring the generated traffic matrices are not only numerically precise but also consistent with realistic network behaviors and topological constraints, while keeping the LLM deployable via parameter-efficient adaptation for proactive and self-adaptive cognitive management.
@@
The remainder of this paper is organized as follows: Section II reviews related works. Section III details the methodology of the proposed LEAD model. Section IV reports our experimental results, including ablation studies. Section V concludes the paper and proposes future research directions.
II Related Works
II-A Network Traffic Prediction Methods
@@
Early methodologies rely heavily on statistical techniques. Historical Average, ARIMA [14], and Kalman filtering [7] are classic examples that model linear dependencies, while Vector Auto-Regression (VAR) [15] extends these to multivariate settings. Despite their efficiency, these models are limited by stationarity assumptions and often struggle to capture the non-linear dynamics of complex networks.
To model non-linear temporal correlations, RNNs, such as LSTM (Long Short-Term Memory) and GRU (Gated Recurrent Unit) [16], are introduced and become foundational baselines. However, these methods treat network traffic streams as isolated sequences, largely neglecting the underlying topological structure and spatial interactions between nodes.
Consequently, Spatio-Temporal Graph Neural Networks (STGNNs) have been widely adopted to jointly model spatial and temporal dependencies. Representative works like DCRNN (Diffusion Convolutional Recurrent Neural Network) [17] and STGCN (Spatial-Temporal Graph Convolutional Network) [18] combine graph convolutions with recurrent or convolutional units to capture directed spatial flows. To mitigate the reliance on predefined graph structures, methods such as Graph WaveNet [19] and AGCRN (Adaptive Graph Convolutional Recurrent Network) [20] propose adaptive graph learning strategies to infer latent correlations automatically. Furthermore, recent studies in the database community have actively explored dynamic and multiscale structure learning to capture evolving dependencies. For instance, METRO [21] introduces a generic framework leveraging multiscale temporal graphs to model variable correlations, while Zhao et al. [22] propose explicitly generating discrete graph structures to adapt to changing time-series relationships. In parallel, researchers have explored advanced representation learning paradigms. For instance, ST-SSL (Spatio-Temporal Self-Supervised Learning) [23] leverages self-supervised learning schemes to enhance robustness against data heterogeneity, while MC-STL (Mask- and Contrast-enhanced Spatio-Temporal Learning) [24] employs multi-channel architectures to disentangle complex trend and seasonality components. More recently, researchers have explored incorporating domain knowledge into STGNNs through attention mechanisms [25].
@@ II-C LLMs
In time-series analysis, for instance, researchers have creatively adapted LLMs by tokenizing numerical sequences. Zhou et al. [32] treat time-series as pseudo-sentences for zero-shot forecasting, while frameworks like Time-LLM [33] and UniTime [34] have demonstrated that a frozen LLM’s internal reasoning capacity can be harnessed for numerical prediction through clever input reprogramming and prompting, effectively bypassing the need for full fine-tuning.
@@ III Methodology
    Encoding: Historical TMs $\textbf{X}_{1:T_{in}}$, where $T_{in}$ denotes the input length of TMs, are encoded into spatio-temporal visual tokens, preserving high-frequency details.
@@
    Generation: These conditions guide a U-Net based Diffusion Model to iteratively denoise Gaussian noise, reconstructing the future TMs $\hat{\textbf{X}}_{{T_{in}+1}:{T_{in}+T_{out}}}$, where $T_{out}$ represents the prediction horizon of TMs.
III-A Data Representation
@@
The transformation’s principal advantage lies in its creation of a representational bridge to convolution neural networks (CNNs). By transposing network data into the visual domain, we leverage the intrinsic capabilities of vision models to identify structurally analogous patterns: intensity gradients become indicators of high-traffic links, while contiguous dark regions reveal network congestion clusters. This approach not only maintains the topological fidelity of the original data but also unlocks sophisticated feature extraction methodologies traditionally applied to visual analysis, thereby facilitating the discovery of spatially distributed patterns within network traffic flows.
III-B Spatio-Temporal Perception via Detail-Preserving Encoding
@@
To aggregate the sequence of visual features $\textbf{H}_{v}^{1:T}$ into a coherent spatio-temporal representation, we introduce an enhanced temporal aggregator. Unlike bidirectional methods including Bi-LSTM (Bidirectional Long Short-Term Memory), network traffic prediction must strictly adhere to causality. We employ a Transformer-based aggregator with a causal mask, ensuring that the representation at step $t$ is based solely on the previous steps.
@@
III-C LLM-Based Reasoning with Adapter
@@
III-D Dual-Conditioning Strategy
@@
III-E Conditional Diffusion Generation
@@
As depicted in Fig. 4, we introduce an Adaptive Group Normalization (AdaGN) mechanism to inject the global context $c_{global}$ into the U-Net backbone. AdaGN dynamically modulates the U-Net’s intermediate feature maps by learning scale and shift parameters derived directly from $c_{global}$, thereby guiding the denoising trajectory with global semantic constraints: $\mathrm{AdaGN}(\mathbf{f},c_{global})=\gamma_{s}(c_{global})\cdot\mathrm{GroupNorm}(\mathbf{f})+\beta_{s}(c_{global})$. Here, $\gamma_{s}$ and $\beta_{s}$ are scale and shift parameters predicted from $c_{global}$ via a shallow Multi-Layer Perceptron. This mechanism effectively shifts the distribution of the generated traffic to match the predicted global congestion level. The sequential condition $c_{seq}$ acts as the Key and Value in the cross-attention layers of the U-Net, while the intermediate feature maps of the U-Net serve as Queries.
@@
During inference, we employ the Denoising Diffusion Implicit Models (DDIM) sampler to accelerate the generation process, allowing us to produce the prediction $\hat{\textbf{X}}_{{T+1}:{T+T_{out}}}$ in fewer steps than the 1000-step denoising process.
@@ IV-B Baselines and Metrics
  • DCRNN & STGCN: DCRNN integrates diffusion convolution with GRU, while STGCN employs graph convolutions and 1D temporal convolutions for faster inference. They both rely on predefined adjacency matrices to capture spatial dependencies [17], [18].
@@
  • ST-SSL & MC-STL: ST-SSL employs self-supervised learning with pretext tasks to learn robust patterns resilient to noise. MC-STL decomposes traffic data into multiple channels to model distinct temporal dynamics separately [23], [24].
@@
  • ViT-LSTM: This method is particularly relevant as it shares the “Traffic-to-Image” paradigm with our approach. It converts TMs into image sequences and utilizes ViT (Vision Transformer) to extract spatial features, followed by LSTM for temporal modeling. By comparing with ViT-LSTM, we can directly validate the superiority of our generative diffusion backbone over traditional discriminative visual models [9].
The primary metrics are pixel-wise root mean squared error (RMSE) and mean absolute error (MAE), computed on the normalized color value of the first channel of RGB images. Given ground truth $X\in\mathbb{R}^{N\times N}$ and prediction $\hat{X}\in\mathbb{R}^{N\times N}$, we have:
@@
IV-C Setup and Implementation Details
@@
To ensure a rigorous evaluation without data leakage, we adopt a strict chronological splitting strategy for both Abilene and GÉANT datasets. The data are divided into training, validation, and testing sets with a ratio of 7:1.5:1.5. Moreover, we employ a sliding window mechanism to generate samples. The model utilizes the historical traffic sequence of length $\bm{T_{in}=24}$ to predict the TMs at the next $\bm{T_{out}}$ time steps.
@@
  • Conditioning: The model integrates global context via FiLM (Feature-wise Linear Modulation) and local details via cross-attention, with a diffusion noise schedule of 1000 steps.
@@
TABLE I: Performance comparison of different models on the Abilene dataset, including RMSE and MAE values.
@@
TABLE II: Performance comparison of different models on the GÉANT dataset, including RMSE and MAE values.
@@
IV-D Main Results and Analysis
The quantitative comparison between our proposed LEAD and nine baselines on the Abilene and GÉANT datasets is presented in Table  II and Table  II respectively, covering prediction horizon from short-term ($T_{out}=1$) to long-term ($T_{out}=20$). Overall, LEAD achieves outstanding performance across both datasets. To provide a more intuitive understanding of the performance divergence across different paradigms, we contrast our model with STGNNs in Fig. 5a-Fig. 5b and with multi-scale methods in Fig. 5c-Fig. 5d.